OpenStudy (anonymous):

the dimension of a rectangle are such that its length is 5 in. more than its width.if the length were doubled and if the width were decreased by 2 in., the area would be increased by 70 in^2 what are the lenghts and widths of the rectangle?

OpenStudy (nubeer):

width =x length = (x+5 ) Area 1=x(x+5) New rectangele Length doubled = 2(x+5) width decrease = x-2 New area is 70 more then the other rectangle so Area 1 + 70 = 2(x+5)(x-2) x(x+5) +70 = 2(x+5)(x-2) now solve this and find x

OpenStudy (anonymous):

i got 20.66666667

OpenStudy (nubeer):

hmm did you got 2 values of x?